OpenStudy (anonymous):

Let u = <4, 3>. Find the unit vector in the direction of u, and write your answer in component form.

OpenStudy (anonymous):

The unit vector of a vector is calculated by dividing the vector by its length. This way, it will keep its direction but have length one.

Parth (parthkohli):



OpenStudy (anonymous):

i got <4/5, 3/5> would that be correct?

Parth (parthkohli):

Yesh!

OpenStudy (anonymous):

ok:) thnk you !

OpenStudy (anonymous):

could u help me on this??.. Two forces with magnitudes of 200 and 100 pounds act on an object at angles of 60° and 170° respectively. Find the direction and magnitude of these forces. Round to two decimal places in all intermediate steps and in your final answer.

OpenStudy (anonymous):

Indeed :) Check the length of the vector if you want to be sure, but you will find: \[\sqrt{\left(\frac{ 4 }{ 5 }\right)^2 + \left( \frac{ 3 }{ 5 } \right)^2} = \sqrt{\frac{ 16 + 9 }{ 25 }} = 1\]

OpenStudy (anonymous):

then 1 is the answer?

OpenStudy (anonymous):

thats to the problem i just did.. how would i do the second one i asked?

OpenStudy (anonymous):

One is the length of the unit vector you just created :)

OpenStudy (anonymous):

ohokay, and do you think u could help me on this??.. Two forces with magnitudes of 200 and 100 pounds act on an object at angles of 60° and 170° respectively. Find the direction and magnitude of these forces. Round to two decimal places in all intermediate steps and in your final answer.